OpenStudy (anonymous):

1. Complete the table below to solve the equation

OpenStudy (anonymous):

1. Complete the table below to solve the equation 2.5x − 10.5 = 64(0.5x). x f(x) = 2.5x − 10.5 g(x) = 64(0.5x) 2  f(x)=  g(x)= 3  f(x)=  g(x)= 4  f(x)=  g(x)= 5  f(x)=  g(x)= 6  f(x)=  g(x)=

OpenStudy (anonymous):

lol idk what comes after 2.5x − 10.5

OpenStudy (mathmale):

You are given 2 different functions, f(x) and g(x). But you don't show what you're supposed to do with them. Add? Subtract? Multiply? Divide? Please go back and ensure that you've copied down all parts of this question.

OpenStudy (anonymous):

this is what it said im sorry there was nothing else

OpenStudy (anonymous):

i actually just don't understand anything about this :( but thank you for trying

OpenStudy (triciaal):

I think you are suppose to use the values of x from 1 to 6 in each function and then see where the result is the same. At that point f(x) = g(x)

OpenStudy (mathmale):

Thanks for the tip, triciaal. Rambler: Please multiply out the right side of the given equation.

OpenStudy (anonymous):

so i multiply what ? the 64 times (0.5)

OpenStudy (mathmale):

64(0.5x) yes

OpenStudy (anonymous):

2.5x − 10.5 = 32x ?

OpenStudy (anonymous):

so then i find what x is?

OpenStudy (mathmale):

Yes. What do you think the next step is?

OpenStudy (anonymous):

x=4?

OpenStudy (mathmale):

Please show ALL steps that you go through.

OpenStudy (anonymous):

i dont think its the right way but i just subrtact 10.5 and then 2.5 get the number than divide by 32?

OpenStudy (anonymous):

to get 4

OpenStudy (anonymous):

2.5(4) − 10.5 = 32(4)???

OpenStudy (mathmale):

This much is correct: 2.5x − 10.5 = 32x You need to "combine like terms." What does that mean?

OpenStudy (anonymous):

take the 2.5 and put it under the 10.5 and subtract?

OpenStudy (anonymous):

or can i not becuase 10 doesnt have an x

OpenStudy (mathmale):

No, because 2.5x and 10.5 are not like terms. Try again.

OpenStudy (anonymous):

but i thought you couldn't combine on opposites of the = sign?

OpenStudy (anonymous):

29.5x - 10.5 ?

OpenStudy (mathmale):

Sign has nothing to do with it here. What's important is that you have "like terms". 1 boy and 1 girl do not equal 2 boys. Why? boy and girl are NOT like terms. Try again.

OpenStudy (anonymous):

so i take the 2.5 and subtract that from 32 because they are alike?

OpenStudy (anonymous):

and get 29.5x

OpenStudy (mathmale):

Yes. More accurately, subtract 2.5x from both sides of the equation.

OpenStudy (anonymous):

so that cancels out 2.5 on the left side

OpenStudy (mathmale):

Now write the equation again, in its latest form.

OpenStudy (anonymous):

29.5x - 10.5 ?

OpenStudy (mathmale):

Where's the = sign? Must keep it. Result must be an equation.

OpenStudy (anonymous):

29.5x - 10.5 = x

OpenStudy (anonymous):

29.5-10.5 =19 so to keep that true would x be 1?

OpenStudy (mathmale):

You should have and get 29.5x on the right side. What does that leave on the left side? To answer your question: there should be no x by itself on the right side. Try again.

OpenStudy (mathmale):

2.5x − 10.5 = 32x

OpenStudy (anonymous):

? what

OpenStudy (anonymous):